OpenStudy (anonymous):

A iron or 4.00m long and 0.500cm2 in cross section stretches 1.00mm when a mass of 225kg is hung from its lower end. Compute the Young's Modulus for the iron

OpenStudy (perl):

what is formula for youngs modulus

OpenStudy (shinalcantara):

\[E=\frac{ PL }{ Ay }\] where E is the young's modulus or the modulus of elasticity P is the applied force L is the length of the material A is the cross-sectional area y is the deformation or elongation P = 225kg(9.81m/s2) = 2,207.25 N L = 4 m = 4000mm A = 0.50cm2(100mm2/1cm2) = 50 mm2 y = 1mm substituting it will be: \[E = \frac{ 2,207.25N(4000mm) }{ 50mm^2(1mm) }\] \[E = 176,580\frac{ N }{ mm^2 }\] \[E = 176,580 MPa\] or expressing it in GPa you'll have \[E = 176.58 GPa\]

OpenStudy (yousir):

y did u block me!!?